You plan to retire 10 years from today.
Your retirement will last for 300 months.
You need $6,000 each month of your retirement.
How much do you have to invest today in an account that earns 6.00% APR (compounded monthly) to exactly pay for your retirement?
Answer #1

Rate = 6% / 12 = 0.5%

Value at year 10 = Annuity *[1 - 1 / (1 + r)n] / r

Value at year 10 = 6,000 *[1 - 1 / (1 + 0.005)300] / 0.005

Value at year 10 = 6,000 * 155.206864

Value at year 10 = $931,241.184

Number of periods = 10 * 12 = 120

Present value = FV / (1 + r)n

Present value = 931,241.184 / (1 + 0.005)120

Present value = 931,241.184 / 1.819397

Present value = $511,840.64

$511,840.64 needs to be invested.